The volume of gastric fluid in your stomach is 100 cm3 and its pH = 2. Calculate the number of H+ ions in your stomach. You need to know that Avogadro’s Number, NA, is 6.02*1023 mol^-1. 


1
Expert's answer
2022-03-16T21:11:03-0400

"[H^+]=10^{-pH}=0.01\ M"


"Number\ of\ H^+=100\ cm^3\times\frac{0.01\ mol}{1000\ cm^3}\times\frac{6.02\times10^{23}\ ions}{1\ mol}=6.02\times10^{20}\ ions"


Answer: "6.02\times10^{20}\ ions"